# references.bib

### v1

@@ Abstract
The computational complexity of polynomial ideals and Gröbner bases has been studied since the 1980s. In recent years the related notions of polynomial subalgebras and SAGBI bases have gained more and more attention in computational algebra, with a view towards effective algorithms. We investigate the computational complexity of the subalgebra membership problem and degree bounds. In particular, we place these problems in the complexity class EXPSPACE and prove PSPACE-completeness for homogeneous algebras. We highlight parallels and differences compared to the settings of ideals and also look at important classes of polynomials such as monomial algebras.
@@ 1 Introduction
Let $\mathbb{K}$ be a field equipped with a suitable encoding over a finite alphabet, for example, the rational numbers $\mathbb{Q}$ or a finite field $\mathbb{F}_{q}$. At the heart of many algorithms in computer algebra are efficient algorithms manipulating polynomial ideals and Gröbner bases, dating back at least to the 1960s [Buchberger2006]. Since the late 1980s, algorithms for computations with subalgebras have been studied [Shannon1988, Robbiano1990], which have applications in toric degenerations and polynomial system solving [betti2025]. A fundamental computational problem for ideals and subalgebras of polynomial rings $\mathbb{K}[{\underline{x}}]=\mathbb{K}[x_{1},\dots,x_{n}]$ is deciding membership. Of special interest are classes C of polynomials such as homogeneous polynomials (Homog), monomials (Mon), or polynomials with bounded degree or number of variables.
@@
Input: $f_{1},\dots,f_{s}\in\textsf{C}$ (or $\mathbb{K}[{\underline{x}}]$),
@@
Input: $f_{1},\dots,f_{s}\in\textsf{C}$ (or $\mathbb{K}[{\underline{x}}]$),
@@
The computational complexity and representation degree bounds of $\textsc{IdealMem}_{\mathbb{K}}$ and its variants have been studied since the 1980s, we give a brief overview in Section 2.3. The complexity of algebras in a more general sense has been studied before [Kozen77], though so far not for the concrete case of polynomial subalgebras. This paper is concerned with filling that gap and proving analogous results for $\textsc{AlgMem}_{\mathbb{K}}$ and an upper bound on the certification degree. In order to obtain reasonable upper bounds on the computational complexity, we assume that arithmetic in the base field $\mathbb{K}$ can be performed efficiently, this is formalized as being well-endowed [Borodin1983]. A main theorem is the PSPACE-completeness of the homogeneous subalgebra membership problem, combining the upper bound Theorem 3.4 and the complementing lower bound Theorem 4.1.
@@
Over a well-endowed field the homogeneous subalgebra membership problem $\emph{{AlgMem}}_{\mathbb{K}}(\textsf{{Homog}})$ is PSPACE-complete.
One of the goals of this paper is to be accessible both to commutative algebraists and computer scientists. For this reason, we introduce the relevant results from algebra and complexity theory on ideals and Gröbner bases in Section 2. In Section 3 we prove various upper bound results on variations of $\textsc{AlgMem}_{\mathbb{K}}$, as well as an upper bound on the certification degree. In Section 4 we present the combinatorial construction of controlled monomial replacement systems which yields a polynomial space lower bound and is of independent interest. In Section 5 we study the case of monomial subalgebras which is still NP-complete, and discuss related questions with SAGBI bases.
@@ 2.1 Ideals, subalgebras and their bases
Let $\mathbb{K}$ be a field, denote by $\mathbb{K}[{\underline{x}}]=\mathbb{K}[x_{1},\dots,x_{n}]$ the polynomial ring in $n$ variables, consisting of finite linear combinations
@@
A monomial order $\prec$ is a total order on the set of monomials $x^{\alpha}$ refining divisibility ($x^{\alpha}\mid x^{\beta}$ implies $x^{\alpha}\preceq x^{\beta}$) and such that $1=x^{0}$ is the minimal element. In the following fix such a monomial order, for example the lexicographic order $\prec_{\text{lex}}$; we may assume $x_{1}>\dots>x_{n}$.
The initial term $\operatorname{in}_{\prec}(f)$ of a nonzero polynomial $f=\sum_{\alpha}c_{\alpha}x^{\alpha}$ is the nonzero term $c_{\alpha}x^{\alpha}$ with the largest monomial. For convenience define $\operatorname{in}_{\prec}(0)=0$. Given an ideal $I$ or a subalgebra $A$, its initial ideal or initial algebra respectively is
@@
A Gröbner basis of an ideal $I$ is a finite subset $G\subseteq I$ such that $\operatorname{in}_{\prec}(I)=\langle\set{\operatorname{in}_{\prec}(g)}{g\in G}\rangle$. On the other hand, a SAGBI basis of a subalgebra $A$ is any subset $S\subseteq A$ such that $\operatorname{in}_{\prec}(A)=\mathbb{K}[\set{\operatorname{in}_{\prec}f}{f\in S}]$. In both cases such bases generate $I$ and $A$ respectively, hence the (historical) name “basis”. A Gröbner basis is reduced if the leading coefficients are $1$ and no term of $g\in G$ is in $\operatorname{in}_{\prec}(G\setminus\{g\})$. The reduced Gröbner basis of an ideal $I$ is unique and has the smallest number of elements and smallest degree among all Gröbner bases of $I$. For more background on Gröbner bases see for example [Cox2016, Chapter 2], [Gathen2017, Chapter 21] or [Kreuzer2000, Chapter 2].
By Hilbert’s basis theorem every ideal can be generated by a finite set of polynomials; in fact one can find such a finite set in any generating set. This implies the existence of (finite) Gröbner bases for any ideal. On the other hand, finite SAGBI bases may not exist, we will come back to this topic in Section 5.1.
Given an ideal $I\subseteq\mathbb{K}[{\underline{x}}]$, any element $f\in\mathbb{K}[{\underline{x}}]$ has a normal form $\operatorname{nf}_{\prec}(f,I)$ against $I$, which is the unique polynomial $r\in f+I$ such that no monomial of $r$ is in $\operatorname{in}_{\prec}(I)$ [Kreuzer2000, Definition 2.4.8]. One can compute $\operatorname{nf}_{\prec}(f,I)$ by dividing $f$ by a Gröbner basis $G$, the remainder is the normal form [Cox2016, Proposition 2.6.1]. This gives a membership test for polynomial ideals: $f\in I$ if and only if $\operatorname{nf}_{\prec}(f,I)=0$.
@@ 2.2 Complexity theory
We use the Turing model of computation, though the details of this formalism are not required to follow this paper. Here we introduce the concepts and computational problems used in the sequel. A classical reference for computational complexity theory is the book by Hopcroft and Ullman [Hopcroft1979], a more modern treatment is the book by Arora and Barak [Arora2009], for a brief introduction with a view towards computer algebra see also [Gathen2017, Section 25.8].
Informally, the complexity of an algorithm is the amount of resources (time or memory) used in the computation as a function in the input length. A decision problem $A$ is the problem of deciding whether a given (well-formed) input $x\in\Sigma^{*}$ has a certain property, i. e. deciding membership in the set $A\subseteq\Sigma^{*}$. Problems in P, PSPACE and EXPSPACE can be solved algorithmically in polynomial time, polynomial space and exponential space respectively, while the answer to problems in NP can be verified in polynomial time provided a certificate. One has the inclusions of complexity classes
@@
A standard problem in NP is Sat, deciding whether a given boolean formula $\varphi(x_{1},\dots,x_{n})$ can be satisfied by some assignment $\{x_{1},\dots,x_{n}\}\to\{\texttt{true},\texttt{false}\}$. We will later need the useful variant 1in3Sat and the SubsetSum problem:
@@
SubsetSum
Input: $a_{1},\dots,a_{n},b\in\mathbb{N}$
Output: Is $\sum_{i\in I}a_{i}=b$ for some $I\subseteq\{1,\dots,n\}$?
@@
The following LBA consists of two states $q_{0},q_{1}$ (plus a halting state), and on input $w\coloneqq\texttt{0}\dots\texttt{0}$ ($n$ zeros) will count in binary to $\texttt{1}\dots\texttt{1}$ and then halt:
@@
Transitions that do not occur (on input $w$) have been left unspecified for simplicity.
A complexity upper bound for a decision problem $A$ and a complexity class C is simply an algorithm solving $A$ and satisfying the resource constraints of C. A lower bound or C-hardness result for $A$ is a proof that the computation of any problem $A^{\prime}\in\textsf{{C}}$ can be reduced to the problem $A$. We use the notion of log-space many-one reductions, in symbols $A^{\prime}\mathrel{\leq^{\textsf{{L}}}_{\text{m}}}A$, this means that there is a map $f$ from inputs of $A^{\prime}$ to inputs of $A$ computable in logarithmic working space, such that $x\in A^{\prime}$ if and only if $f(x)\in A$. For example, SubsetSum and 1in3Sat are NP-complete and LBAhalt is PSPACE-complete.
In the problems $\textsc{IdealMem}_{\mathbb{K}}$ and $\textsc{AlgMem}_{\mathbb{K}}$ the input consists of encoded polynomials over $\mathbb{K}$, for example as a string of characters, so the input length is bounded below by the number of terms, the number of variables occurring and (depending on the encoding) the coefficients and the (logarithm of the) degree. All our complexity results will hold with respect to both dense and sparse encodings and binary or unary monomial representation (with the exception of Theorem 5.2), as long as we make the reasonable assumption that the field $\mathbb{K}$ is well-endowed [Borodin1983]. Commonly used fields in computer algebra such as number fields and finite fields are well-endowed. For this reason, we will not elaborate further on the encoding.
2.3 Brief summary of the ideal world
@@
Note that this uses the convention that in an enumeration problem we only consider the working space and not the cells on the output tape. For example, one can enumerate the binary numbers $10,1,\dots,2^{n}-$ 1 with only about $n$ bits of working memory. Similarly, here the length of a certificate may be doubly exponential in the input length, this is unavoidable in the worst case [Mayr1982].
Regarding representation degree bounds are in general doubly exponential in the number of variables, which is asymptotically optimal. We also note the special case of complete intersections for later reference. For a definition of complete intersections see e. g. [Cox2016, Exercise 9.4.8] or [Kreuzer2000, Definition 3.2.23].
@@
Both from a theoretical and computational point of view, effective degree bounds on Gröbner bases are also important. We recall the classical Dubé bound as well as a dimension-refined version.
@@
    (Dubé [Dube1990]) The degree of the reduced Gröbner basis GB of $I$, i. e. the largest degree of an element of GB, is bounded above by
    $\deg\textit{GB}\leq 2\left(\frac{d_{1}^{2}}{2}+d_{1}\right)^{2^{n-1}}.$
@@
    (Mayr & Ritscher [Mayr2013]) If $\dim\mathbb{K}[{\underline{x}}]/I=r$, then the degree is bounded by
@@ 3 Upper bounds on subalgebra membership
To prove better complexity upper bounds for special classes of polynomials we outline part of Kühnle & Mayr’s upper bound constructions on normal forms [Kühnle1996].
@@
For a fixed subalgebra $A=\mathbb{K}[f_{1},\dots,f_{s}]\subseteq\mathbb{K}[{\underline{x}}]$ over a well-endowed field, the membership problem is in PSPACE (with respect to the input length of $g$). In fact, bounding the number of variables $n$ already implies $\emph{{AlgMem}}_{\mathbb{K}}(\mathbb{K}[x_{1},\dots,x_{n}])\in\textsf{{PSPACE}}$.
@@
If $g\in\mathbb{K}[f_{1},\dots,f_{s}]\subseteq\mathbb{K}[{\underline{x}}]$, then there exists a certificate $h\in\mathbb{K}[t_{1},\dots,t_{s}]$ with $g=h(f_{1},\dots,f_{s})$ of degree
$\deg h\leq\deg(g)+\left(\left(\textstyle\frac{1}{2}d^{2s^{2}}+d\right)^{2^{n}}+1\right)^{(n+s)^{2}+1}\deg(g)^{n+s}.$
@@
As a consequence of the grading, we see that a certificate of lowest degree (if it exists) has degree at most $\deg g$. Thus either $\deg\operatorname{nf}_{\prec}(h,J)>\deg d$, in which case $g\notin K[f_{1},\dots,f_{s}]$, or $\deg\operatorname{nf}_{\prec}(h,J)\leq\deg d$ and $g\in\mathbb{K}[f_{1},\dots,f_{s}]$ if and only if $\operatorname{nf}_{\prec}(h,J)\in K[{\underline{t}}]$. This leads to a variation of the algorithm of Theorem 3.1 except we only attempt to compute the normal form up to degree $\deg g$. The complexity analysis is analogous to Theorem 3.2, noting that here a representation degree bound is $D\coloneqq\deg g+d^{s}$. ∎
@@
4 Lower bounds for homogeneous subalgebras
In this section, we prove a matching lower bound for the homogeneous subalgebra membership problem, similar to the construction for homogeneous ideals. We will describe a combinatorial problem we call controlled monomial replacement systems (CMRS), which is PSPACE-hard, and which we can embed into the homogeneous subalgebra membership problem. The hardness results for $\textsc{IdealMem}_{K}$ and $\textsc{IdealMem}_{K}(\textsf{{Homog}})$ were constructed by embedding the word problem for commutative semigroups resp. its homogeneous variant into binomial ideals [Mayr1982, Mayr1997]. Our construction is similar but a bit trickier since the embedding of CMRS into subalgebras is less natural.
Consider a set of variables ${\underline{x}}=\mathcal{C}\mathbin{\dot{\cup}}\mathcal{P}$ partitioned into $\mathcal{P}$adding variables and $\mathcal{C}$ontrol variables, $\mathcal{C}\neq\emptyset$. Given a set of binomial replacement rules =$\mathcal{R}=\{x^{\alpha_{i}}-x^{\beta_{i}}\}_{i=1}^{r}$ we define an equivalence relation $\equiv_{\mathcal{R}}$ on the set of monomials $\textsf{{Mon}}({\underline{x}})$ generated by
$x^{\gamma}x^{\alpha}\equiv_{\mathcal{R}}x^{\gamma}x^{\beta}\qquad x^{\alpha}-x^{\beta}\in\mathcal{R},\quad x^{\gamma}\in\textsf{{Mon}}(\mathcal{P}).$ (1)
The application of a single rule $R\in\mathcal{R}$ “from left to right” will be indicated by $m\equiv^{R}m^{\prime}$. The application in the other direction is $m^{\prime}\equiv^{-R}m$, which we call a reverse step (assuming $\mathcal{R}\cap-\mathcal{R}=\emptyset$).
We call a polynomial $f$ controlled with respect to a partition $\mathcal{C}=\mathcal{C}_{1}\mathbin{\dot{\cup}}\dots\mathbin{\dot{\cup}}\mathcal{C}_{m}$ if $f$ is homogeneous of degree one in each $\mathcal{C}_{i}$. This means that each monomial in $f$ contains exactly one variable from each $\mathcal{C}_{i}$, and otherwise only padding variables.
[…]
Input: Variables ${\underline{x}}=\mathcal{C}_{1}\mathbin{\dot{\cup}}\dots\mathbin{\dot{\cup}}\mathcal{C}_{m}\mathbin{\dot{\cup}}\mathcal{P}$,
(homogeneous) controlled replacement rules =$\mathcal{R}=\{x^{\alpha_{i}}-x^{\beta_{i}}\}_{i=1}^{r}$,
controlled monomials $x^{\alpha},x^{\beta}\in\textsf{{Mon}}({\underline{x}})$
Output: Is $x^{\alpha}\equiv_{\mathcal{R}}x^{\beta}$?
[…]
{lemma}
[…]

### v2

@@ Abstract
The computational complexity of polynomial ideals and Gröbner bases has been studied since the 1980s. In recent years, the related notions of polynomial subalgebras and SAGBI bases have gained more and more attention in computational algebra, with a view towards effective algorithms. We investigate the computational complexity of the subalgebra membership problem and degree bounds. In particular, we show completeness for the complexity class EXPSPACE and prove PSPACE-completeness for homogeneous algebras. We highlight parallels and differences compared to the settings of ideals, and also look at important classes of polynomials such as monomial algebras.
@@ 1 Introduction
Let $\mathbb{K}$ be a field equipped with a suitable encoding over a finite alphabet, for example, the rational numbers $\mathbb{Q}$ or a finite field $\mathbb{F}_{q}$. At the heart of many algorithms in computer algebra lie efficient algorithms manipulating polynomial ideals and Gröbner bases, dating back at least to the 1960s [Buchberger2006]. Since the late 1980s, algorithms for computations with subalgebras have been studied [Shannon1988, Robbiano1990], which have applications in toric degenerations and polynomial system solving [betti2025]. A fundamental computational problem for ideals and subalgebras of polynomial rings $\mathbb{K}[{\underline{x}}]=\mathbb{K}[x_{1},\dots,x_{n}]$ is deciding membership. Of special interest are classes C of polynomials such as homogeneous polynomials (Homog), monomials (Mon), or polynomials with bounded degree or number of variables.
@@
Input: $f_{1},\dots,f_{s}\in\textsf{C}$ (or $\mathbb{K}[{\underline{x}}]$);
@@
Input: $f_{1},\dots,f_{s}\in\textsf{C}$ (or $\mathbb{K}[{\underline{x}}]$);
@@
The computational complexity and representation degree bounds of $\textsc{IdealMem}_{\mathbb{K}}$ and its variants have been studied since the 1980s, we give a brief overview in Section 2.3. The complexity of algebras in a more general sense has been studied before [Kozen77], though so far not for the concrete case of polynomial subalgebras. This paper is concerned with filling this gap and proving analogous results for $\textsc{AlgMem}_{\mathbb{K}}$ and bounds on the certification degree. In order to obtain reasonable upper bounds on the computational complexity, we assume that arithmetic in the base field $\mathbb{K}$ can be performed efficiently; this is formalized as being well-endowed [Borodin1983]. A main theorem is the EXPSPACE- and PSPACE-completeness of the subalgebra membership problem and the homogeneous variant, combining the upper bounds Theorem 3.1 and 3.4 and the complementing lower bound Theorem 4.2.
@@
Over a well-endowed field $\mathbb{K}$, the subalgebra membership problem $\emph{{AlgMem}}_{\mathbb{K}}$ is EXPSPACE-complete and the homogeneous version $\emph{{AlgMem}}_{\mathbb{K}}(\textsf{{Homog}})$ is PSPACE-complete.
One of the goals of this paper is to be accessible both to commutative algebraists and computer scientists. For this reason, we introduce the relevant results from algebra and complexity theory on ideals and Gröbner bases in Section 2. In Section 3 we prove various upper bound results on variations of $\textsc{AlgMem}_{\mathbb{K}}$, as well as an upper bound on the certification degree. In Section 4 we present a complexity-theoretic reduction from $\textsc{IdealMem}_{\mathbb{K}}$ to $\textsc{AlgMem}_{\mathbb{K}}$, proving matching lower bounds. In Section 5 we study the case of monomial subalgebras, which is still NP-complete, and discuss related questions with SAGBI bases.
@@ 2.1 Ideals, subalgebras and their bases
Let $\mathbb{K}$ be a field, denote by $\mathbb{K}[{\underline{x}}]\coloneqq\mathbb{K}[x_{1},\dots,x_{n}]$ the polynomial ring in $n$ variables, consisting of finite linear combinations
@@
Definition 2.3.
A monomial order $\prec$ is a total order on the set of monomials $x^{\alpha}$ stable under multiplication ($x^{\alpha}\preceq x^{\beta}$ implies $x^{\alpha}x^{\gamma}\preceq x^{\beta}x^{\gamma}$) and such that $1=x^{0}$ is the minimal element. The initial term $\operatorname{in}_{\prec}(f)$ of a nonzero polynomial $f=\sum_{\alpha}c_{\alpha}x^{\alpha}$ is the nonzero term $c_{\alpha}x^{\alpha}$ with the largest monomial. For convenience define $\operatorname{in}_{\prec}(0)=0$. Given an ideal $I$ or a subalgebra $A$, its initial ideal or initial algebra, respectively, is
@@
In the following fix such a monomial order, for example, the lexicographic order $\prec_{\text{lex}}$; we may assume $x_{1}>\dots>x_{n}$.
Definition 2.4.
A Gröbner basis of an ideal $I$ is a finite subset $G\subseteq I$ such that $\operatorname{in}_{\prec}(I)=\langle\set{\operatorname{in}_{\prec}(g)}{g\in G}\rangle$. A Gröbner basis is reduced if the leading coefficients are $1$ and no term of $g\in G$ is in $\operatorname{in}_{\prec}(G\setminus\{g\})$.
On the other hand, a SAGBI basis of a subalgebra $A$ is any subset $S\subseteq A$ such that $\operatorname{in}_{\prec}(A)=\mathbb{K}[\set{\operatorname{in}_{\prec}(f)}{f\in S}]$.
In both cases, such bases generate $I$ and $A$ respectively, hence the (historical) name “basis”. The reduced Gröbner basis of an ideal $I$ is unique and has the smallest number of elements and smallest degree among all Gröbner bases of $I$. For more background on Gröbner bases see for example [Cox2016, Chapter 2], [Gathen2017, Chapter 21] or [Kreuzer2000, Chapter 2].
By Hilbert’s basis theorem, every ideal can be generated by a finite set of polynomials; in fact one can find such a finite set in any generating set. This implies the existence of (finite) Gröbner bases for any ideal. On the other hand, finite SAGBI bases may not exist; we will come back to this topic in Section 5.1.
Definition 2.5.
Given an ideal $I\subseteq\mathbb{K}[{\underline{x}}]$, the normal form $\operatorname{nf}_{\prec}(f,I)$ of an element $f\in\mathbb{K}[{\underline{x}}]$ against $I$ is the unique polynomial $r\in f+I$ such that no monomial of $r$ is in $\operatorname{in}_{\prec}(I)$ [Kreuzer2000, Definition 2.4.8].
One can compute $\operatorname{nf}_{\prec}(f,I)$ by dividing $f$ by a Gröbner basis $G$, the remainder is the normal form [Cox2016, Proposition 2.6.1]. This gives a membership test for polynomial ideals: $f\in I$ if and only if $\operatorname{nf}_{\prec}(f,I)=0$.
@@ 2.2 Complexity theory
We use the Turing model of computation, though the details of this formalism are not required to follow this paper. Here we introduce the concepts and computational problems used in the sequel. A classical reference for computational complexity theory is the book by Hopcroft and Ullman [Hopcroft1979], a more modern treatment is the book by Arora and Barak [Arora2009], for a brief introduction with a view towards computer algebra, see also [Gathen2017, Section 25.8].
Informally, the complexity of an algorithm is the amount of resources (time or memory) used in the computation as a function of the input length.
Definition 2.6.
A decision problem $A$ is the problem of deciding whether a given input $x\in\Sigma^{*}\coloneqq\{\text{words over the alphabet }\Sigma\}$ has a certain property, i. e. deciding membership in the set $A\subseteq\Sigma^{*}$. Problems in P, PSPACE, and EXPSPACE can be solved algorithmically in polynomial time, polynomial space, and exponential space, respectively, while the answer to problems in NP can be verified in polynomial time provided a certificate.
One has the inclusions of complexity classes
@@
A standard problem in NP is Sat, deciding whether a given boolean formula $\varphi(x_{1},\dots,x_{n})$ can be satisfied by some assignment $\{x_{1},\dots,x_{n}\}\to\{\texttt{true},\texttt{false}\}$. We will later need the useful variant 1in3Sat (compare [Schaefer78], also known as positive/monotone 1in3Sat) and the UnboundedSubsetSum problem:
@@
UnboundedSubsetSum
Input: $a_{1},\dots,a_{s};b\in\mathbb{N}$
Output: Is $\sum_{i=1}^{s}c_{i}a_{i}=b$ for some $c_{1},\dots,c_{s}\in\mathbb{N}$?
@@
The following table shows the transition function of a LBA consists of two states $q_{0},q_{1}$ (plus a halting state), and on input $w\coloneqq\texttt{0}\dots\texttt{0}$ ($n$ zeros) will count in binary to $\texttt{1}\dots\texttt{1}$ and then halt:
@@
Transitions that do not occur (on input $w$) have been left unspecified for simplicity. An example of the transition sequence can be found at
mathrepo.mis.mpg.de/ComplexityOfSubalgebras/configurations.html
Definition 2.8.
A complexity upper bound for a decision problem $A$ and a complexity class C is an algorithm solving $A$ and satisfying the resource constraints of C.
A lower bound or C-hardness result for $A$ is a proof that the computation of any problem $A^{\prime}\in\textsf{{C}}$ can be reduced to the problem $A$. We use the notion of log-space many-one reductions:
Definition 2.9.
For two sets $A\subseteq\Sigma^{*}$, $A^{\prime}\subseteq(\Sigma^{\prime})^{*}$, the problem $A^{\prime}$ is log-space many-one reducible to $A$, in symbols $A^{\prime}\mathrel{\leq^{\textsf{{L}}}_{\textup{m}}}A$, if there is a map $f\colon(\Sigma^{\prime})^{*}\to\Sigma^{*}$ computable in logarithmic working space, such that $x\in A^{\prime}$ if and only if $f(x)\in A$.
A problem $A$ is C-hard, if $A^{\prime}\mathrel{\leq^{\textsf{{L}}}_{\textup{m}}}A$ for all $A^{\prime}$ in C. It is C-complete if it is also in C.
For example, Sat, 1in3Sat, and UnboundedSubsetSum are NP-complete, and LBAhalt is PSPACE-complete.
In the problems $\textsc{IdealMem}_{\mathbb{K}}$ and $\textsc{AlgMem}_{\mathbb{K}}$, the input consists of encoded polynomials over $\mathbb{K}$, for example as a string of characters. Choosing a reasonable representation, the input length is bounded below by the number of terms, the number of variables occurring, and, depending on the encoding, the coefficients and the (logarithm of the) degree.
Definition 2.10.
The encoding of a polynomial $f=\sum_{|\alpha|\leq d}c_{\alpha}x^{\alpha}$ is dense if it lists all monomials with their coefficients $(\alpha,c_{\alpha})$ up to its degree, and sparse if it only lists those terms with non-zero coeffient. The representation of a monomial $x^{\alpha}\mathbin{\hat{=}}(\alpha_{1},\dots,\alpha_{n})\in\mathbb{N}^{n}$ is in binary if each $\alpha_{i}$ is encoded in binary, and unary if it is encoded in unary.
Example 2.11.
The term $f=42x_{1}^{3}x_{3}^{6}\in\mathbb{Q}[x_{1},x_{2},x_{3}]$ could be encoded with binary exponents as $((11,0,110),\texttt{enc}(42))$, and with unary as $((111,0,111111),\texttt{enc}(42))$. In dense encoding, a representation of $f$ would have to list all $\binom{9+3}{3}=220$ monomials up to degree $9$ (which most have coefficient 00), while a sparse encoding only has to list that one term.
All our complexity results will hold with respect to both dense and sparse encodings and binary or unary monomial representation (with the exception of Theorem 5.2), as long as we make the reasonable assumption that the field $\mathbb{K}$ is well-endowed [Borodin1983]. Commonly used fields in computer algebra such as number fields and finite fields are well-endowed. For this reason, we will not elaborate further on the encoding.
2.3 The ideal world: brief summary of results
@@
Note that this uses the convention that in a computational problem with output, we only consider the working space and not the space needed to store the entire output. For example, one can enumerate the binary numbers $10,1,\dots,2^{n}-$ 1 with only $\mathcal{O}(n)$ bits of working memory. Similarly, here the number of terms in a representation $(h_{i})_{i}$ may be doubly exponential in the input length; this is unavoidable in the worst case [Mayr1982].
Representation degree bounds are, in general, doubly exponential in the number of variables, which is asymptotically optimal. We also note the special case of complete intersections for later reference. For a definition of complete intersections see e. g. [Cox2016, Exercise 9.4.8] or [Kreuzer2000, Definition 3.2.23].
@@
Both from a theoretical and computational point of view, effective degree bounds on Gröbner bases are important as well. We recall the classical Dubé bound as well as a dimension-refined version. Here, the dimension of an ideal $I\subseteq\mathbb{K}[{\underline{x}}]$ is the (Krull) dimension of the quotient ring $\mathbb{K}[{\underline{x}}]/I$, see [Cox2016, Chapter 9] for other characterizations of the dimension.
@@
    (Dubé [Dube1990]) The degree of the reduced Gröbner basis GB of $I$, i. e. the largest degree of an element of GB, is bounded by
    $\deg\textit{GB}\leq 2\left(\frac{1}{2}d_{1}^{2}+d_{1}\right)^{2^{n-1}}.$
@@
    (Mayr & Ritscher [Mayr2013]) If $I$ has dimension $r$, then the degree is bounded by
@@ 3 Upper bounds on subalgebra membership
To prove degree bounds and better complexity upper bounds for special classes of polynomials, we outline part of Kühnle & Mayr’s upper bound constructions on normal forms [Kühnle1996].
@@
For a fixed subalgebra $A=\mathbb{K}[f_{1},\dots,f_{s}]\subseteq\mathbb{K}[{\underline{x}}]$ over a well-endowed field, the membership problem is in PSPACE (with respect to the input length of $g$). In fact, for a fixed number of variables $n$, we have $\emph{{AlgMem}}_{\mathbb{K}}(\mathbb{K}[x_{1},\dots,x_{n}])\in\textsf{{PSPACE}}$.
@@
If $g\in\mathbb{K}[f_{1},\dots,f_{s}]\subseteq\mathbb{K}[{\underline{x}}]$, then there exists a certificate $p\in\mathbb{K}[t_{1},\dots,t_{s}]$ with $g=p(f_{1},\dots,f_{s})$ of degree
$\deg p\leq\deg(g)+\left(\left(\textstyle\frac{1}{2}d^{2s^{2}}+d\right)^{2^{n}}+1\right)^{(n+s)^{2}+1}\deg(g)^{n+s}.$
@@
As a consequence of the grading, we see that a certificate $p$ of lowest degree (if it exists) has degree at most $\deg g$. Thus either $\deg\operatorname{nf}_{\prec}(g,J)>\deg g$, in which case $g\notin K[f_{1},\dots,f_{s}]$, or $\deg\operatorname{nf}_{\prec}(g,J)\leq\deg g$ and $g\in\mathbb{K}[f_{1},\dots,f_{s}]$ if and only if $\operatorname{nf}_{\prec}(g,J)\in K[{\underline{t}}]$. This leads to a variation of the algorithm of Theorem 3.1 except we only attempt to compute the normal form up to degree $\deg g$. The complexity analysis is analogous to Theorem 3.2, noting that $\deg\operatorname{nf}_{\prec}(g,J)\leq\deg g$, hence a representation degree bound here is $D\coloneqq\deg g+d^{s}$. ∎
@@
4 Lower bounds on subalgebra membership
In this section, we prove matching complexity lower bounds for the subalgebra membership problem and its homogeneous variant, relating them to the construction for homogeneous ideals. The hardness results for $\textsc{IdealMem}_{\mathbb{K}}$ and $\textsc{IdealMem}_{\mathbb{K}}(\textsf{{Homog}})$ were constructed by embedding the word problem for commutative semigroups resp. its homogeneous variant into binomial ideals [Mayr1982, Mayr1997]. We now describe how to embed (binomial) ideal membership into subalgebra membership.
Theorem 4.1.
Let $f_{1},\dots,f_{s},g\in\mathbb{K}[x_{1},\dots,x_{n}]$ and let $t$ be a new variable. The following are equivalent:
  1. (a)
    $g\in\langle f_{1},\dots,f_{s}\rangle\subseteq\mathbb{K}[{\underline{x}}]$;
  2. (b)
    $tg\in\mathbb{K}[tf_{1},\dots,tf_{s},x_{1},\dots,x_{s}]$.